### v1

# Instruct-Tuning Pretrained Causal Language Models for Ancient Greek Papyrology and Epigraphy

Abstract
This article presents an experiment in fine-tuning a pretrained causal language model (Meta’s Llama 3.1 8B Instruct) for aiding in three fundamental tasks of philological research: chronological and geographic attribution as well as text restoration in ancient Greek inscriptions and documentary papyri. Using a prompt-based instruct approach, the fine-tuned models surpass the state of the art in key metrics. For inscriptions, the models achieve a lower average character error rate (CER) of 22.5% (vs. 26.3%), while closely matching top-1 accuracy (60.9% vs. 61.8%) and top-20 accuracy (77.5% vs. 78.3%) for sequences up to 10 characters. They also provide a practical advantage by ignoring spaces during reconstruction, aligning better with the scriptio continua format of ancient written artifacts. In geographic attribution, the model outperforms previous benchmarks with a top-1 accuracy of 75.0% (vs. 70.8%) and a top-3 accuracy of 83.7% (vs. 82.1%). For dating, it achieves an average deviation of 26.2 years (vs. 29.3) and a median deviation of 1 year (vs. 3) from the actual date range. The models also set new baselines for documentary papyri, with a CER of 16.3%, a top-1 accuracy of 71.3%, and top-20 of 85.0% in text reconstruction; a top-1 accuracy of 66.4% and top-3 of 79.9% in geographic attribution; and, in chronological attribution, a deviation of 21.7 years from the actual termini post/ante quem, with a median deviation of 0 years.
Keywords papyrology  $\cdot$ epigraphy  $\cdot$ philology  $\cdot$ deep learning  $\cdot$ large language models  $\cdot$ instruct models
1 Introduction
A significant number of documentary texts in ancient Greek—such as announcements, laws, receipts, contracts, and letters—have survived from the Archaic period through to the early Byzantine era. These texts were inscribed or written on various materials, often on durable stone blocks or sheets of papyrus, which have survived particularly well in drier climates.
A key task of philological research is to assign dates to these documents and attribute them to their places of origin. Insights can typically be drawn from the provenance of the artifact in question, its physical features and writing style, but typically the analysis also relies on clues from the textual content, such as the information conveyed or specific linguistic features. Another important task is the reconstruction of damaged or missing letters in the fragments, which are frequently marred by gaps, abrasions, and scribal errors. Grammatical and stylistic criteria, as well as parallels from similar documents—now readily searchable through databases like www.papyri.info and inscriptions.packhum.org—can be invoked to support or challenge specific conjectures or emendations. However, the conception of such suggestions has traditionally relied on the linguistic creativity and inventiveness of individual scholars, and on their personal expertise acquired over many years of reading and working with similar ancient texts. Conjectural “higher" criticism is performed ope ingenii (“with the aid of one’s ingenuity”), as textual critics put it. It is a branch of philological research that often seems to align more closely with the realm of art than with that of science, or at least with a scientific field where, to borrow a phrase from T.S. Eliot [1, p. 103], “there is no method except to be very intelligent.”
This traditional division within textual criticism—between the objective, technical analysis of manuscript data and the subjective, intuitive insight of an astute critic—is undergoing a significant transformation. Machine learning has the potential to contribute across all areas of philology, including those traditionally characterized by interpretation, guesswork, and creativity. Language models can be pretrained on the entire surviving textual record of ancient civilizations and fine-tuned to parse, classify, and analyze various aspects of words, phrases, and texts in these languages. They can be applied to explore relationships between manuscripts, identify prosodic, rhetorical, and narrative patterns, detect intertextual relationships, recognize themes, analyze sentiments, summarize ideas, and translate texts (see review by Sommerschield and colleagues [2]). More immediately relevant to our purposes, they can be trained to date and geographically attribute texts, detect errors, generate corrections, and fill in missing words and phrases ([3, 4, 5, 6, 7]).
In the field of Greek documentary texts, Assal and colleagues [4] remain the leading pioneers. They trained a transformer model with a sparse attention mechanism (BigBird), combining word- and character-level tokenization, to date, attribute geographically, and reconstruct damaged sequences in ancient Greek inscriptions. On their test set of 7,811 inscriptions, they reported the following results:
  • Text Restoration: The model achieved a top-1 accuracy of 61.8% for restoring sequences of 1 to 10 characters, compared to 25.3% for the human experts they recruited. The character error rate (CER) for the model was 26.3%, while the CER for the philologists was 59.6%. Additionally, 78.3% of the correct cases were found within the top 20 results.
  • Geographical Attribution: The model reached a top-1 accuracy of 70.8%, with the correct location appearing within the top 3 in 82.1% of cases.
  • Chronological Attribution: The model’s predictions were, on average, 29.3 years away from the actual dating span (termini post and ante quem), with a median error of only 3 years.
The modest aim of this experiment was to advance research in machine learning and textual criticism in three respects:
First, the model published by Assael and colleagues [4] counts spaces as characters in the restoration task. For example, the sequence ϰαι ο λογ – – – – – – ος τον ϑεον would not be decoded as ϰαι ο λογος ην προς τον ϑεον. Instead, the input would need to be ϰαι ο λογ – – – – – – – – ος τον ϑεον, forcing us to guess beforehand that damaged characters 3 and 6 are empty spaces. In ancient inscriptions and papyri—typically written in scriptio continua or at least without evenly marked spaces—we often know the approximate number of missing letters but not the combined number of missing characters and word boundaries. This study aims to develop a model that tracks only characters, better reflecting the real-world situations philologists face when working with damaged textual artifacts.
Second, given the rapid advancements in deep learning and Natural Language Processing over the last two years, it is worth exploring how newer, larger pretrained causal language models—such as Meta’s LLaMA, Google’s Gemma, Microsoft’s Phi, or OpenAI’s GPT-4o—can assist textual critics. One promising approach could involve enabling masked language modeling in these models and retraining them to fill in missing sequences of text within a bidirectional context (see the first step of the recipe developed by [8], LLM2Vec). However, in this preliminary study, I have opted for a much simpler approach that is immediately applicable and open to experimentation with larger and newer models in the future: using a straightforward instruct template prompting the model to generate the date, location, or reconstruction at the end of a sequence.
Third, I aimed to expand the application of deep learning for textual criticism of ancient Greek documentary texts to include papyri in addition to inscriptions. While Pavlopoulos and colleagues [9] developed a regression model that dates Greek papyri with an average error of 54 years, there has otherwise been little strictly text-based deep-learning research on documentary papyri.
2 Method
Date spans, geographic attributions, and texts for the ancient Greek inscriptions were sourced from the Packard Humanities Institute database (inscriptions.packhum.org/) using the text processing pipeline developed and shared by Assael and colleagues [4] (github.com/sommerschield/iphi). The texts of Greek documentary papyri were obtained from the Duke Databank of Documentary Papyri (DDbDP), with their chronological and geographical metadata sourced from the Heidelberger Gesamtverzeichnis der griechischen Papyrusurkunden Ägyptens (HGV), both accessible via https://papyri.info.
To facilitate comparison with the results reported by Assael and colleagues [4], the datasets were cleaned of diacritics. Commas were removed and all other punctuation marks converted to a high dot “·”. The data was then split into a train (95%) and test sets (5%).
Dates for both papyri and inscriptions were formatted as exact integers corresponding to the midpoint of their respective date spans. For dates with only a terminus post quem, they were set 25 years after the given date, while those with only a terminus ante quem were set 25 years before. Place name attributions for the papyri were normalized semi-manually due to inconsistencies across database entries. Although more detailed processing would be beneficial in the future, the current approach – admittedly rough (see https://github.com/ericu9500/PapyriAndInscriptions/train_data/data/places_and_dates.tsv) – was deemed sufficient for the experimental proof-of-concept purposes of this study.
All text editions were included in two versions: one in which philologists’ reconstructions of lost or uncertain characters were integrated, and another where these characters were replaced with “-”. Individual missing characters were represented by a single “-”, while longer stretches of missing text were indicated by 10 consecutive “-”.
Text masks were generated before tokenization by randomly selecting spans of 3 to 20 characters, ensuring that no more than 50% of a sequence of intact characters was masked. Longer spans of intact characters were favored by assigning weights proportional to the square of their lengths, increasing the probability of selecting longer spans for masking. These spans were replaced with a simple placeholder in the format “[6 letters missing]”, where the count excluded spaces and punctuation marks within the masked text segment.
The training data was formatted using the same chat template that the instruct model had been pretuned with, utilizing one of three system prompts: “Date this papyrus fragment/inscription to an exact year!”, “Assign this papyrus fragment/inscription to an exact place!”, or “Reconstruct the missing letters in this papyrus fragment/inscription!”. The user input consisted of a version of the text, while the assistant’s response provided the corresponding date, place, or masked sequence of characters, including spaces and punctuation marks. The text was tokenized using the model’s TikToken-based tokenizer.
The model, LLaMA 3.1 8B-Instruct [10], was fine-tuned on 4 A100 GPUs with 40 GB and batch size of 8 or 80 GB and batch size of 20 depending on availability at the time, utilizing the toolkit Torchtune (github.com/pytorch/torchtune). In an initial calibration round, exploratory fine-tuning was conducted with three different data combinations to identify the best approach: one model trained solely on papyri (4 epochs, approximately 80 hours), another solely on inscriptions (4 epochs, approximately 120 hours), and a third on both inscriptions and papyri (3 epochs, approximately 132 hours). For each data combination, two versions were trained: one using the entire sequence of system, user, and assistant messages, and another with the prompt masked out for loss computation.
Subsequently, the best-performing models were further trained separately for the three tasks until their performance plateaued on the test set. During this stage, entries shorter than 75 tokens or longer than 847 tokens were filtered out, and data augmentation was applied by randomly altering the order of sentences and adding noise through the replacement of 5%, 10%, 15%, 20%, and 25% of preserved characters with “-”. Also, masked sequences of 1–20 (and not only 3-20) characters were included. In the final stage, the fine-tuned models were re-merged with the base model using the TIES method developed by Yadav and colleagues [11] with the toolkit Mergekit (github.com/arcee-ai/mergekit).
The models were tested using the same metrics proposed by Assael and colleagues [4]: CER (Character Error Rate), top-1 and top-20 accuracy for normalized reconstructions (spaces, punctuation marks and numerals ignored), top-1 and top-3 accuracy for geographical attributions, and average and median deviation from the date ranges provided in the database. Entries shorter than 90 characters were excluded. For inscriptions, the test was conducted on 7,811 samples for restoration, 2,870 samples for geographical attribution, and 1,856 samples for dating. For papyri, the test was conducted on 7,811 samples for restoration, 1,990 samples for geographical attribution, and 2,295 samples for dating. Dates, places, and reconstructions were evaluated based on top-1, top-3, and top-20 rankings, respectively, out of 45 beams. For reconstructions, a repetition penalty of 1.2 was applied to encourage varied predictions.
3 Results
3.1 Calibration
For the reconstruction task, training with the prompt masked out was clearly superior, resulting in a difference of 8% CER for the inscriptions and 7% for the papyri after 4 epochs, all other factors being equal. The transfer effect of training the same model on both papyri and inscriptions was slightly negative, with a 5% difference for inscriptions and 0.1% difference for papyri after 3 epochs.
For the geographical attribution task, training with the prompt masked out was also superior, yielding a 10.8% difference in top-1 accuracy for inscriptions and a 4.9% difference for papyri after 4 epochs. The transfer effect of training the same model on both papyri and inscriptions was negative, with a significant 39.2% difference for inscriptions and a smaller 1.8% difference for papyri after 3 epochs.
For the dating task, again, training with the prompt masked out led to better performance, with an average difference of 16 years for inscriptions and 3 years for papyri after 4 epochs. The transfer effect of training the same model on both papyri and inscriptions was slightly positive for papyri (-2 years) but negative for inscriptions (0.3 years), after 3 epochs.
The top models in this round were as follows:
  • Reconstruction: Three epochs for inscriptions (CER 28.1%, top-1 accuracy 52.8%, and top-20 accuracy 68.4%) and four epochs for papyri (CER 21.0%, top-1 accuracy 64.9%, and top-20 accuracy 78.3%).
  • Geographical Attribution: Three epochs for inscriptions (top-1 accuracy 71.0%, top-3 accuracy 81.4%) and four epochs for papyri (top-1 accuracy 62.3%, top-3 accuracy 77.9%).
  • Dating: Three epochs for inscriptions (average distance 40.5 years, median distance 5 years) and three epochs (trained on both papyri and inscriptions) for papyri (average distance 22.11 years, median distance 0 years).
3.1.1 Continued training
The calibration round provided several key insights: it is beneficial to mask out the prompt; the effects of transfer learning between papyri and inscriptions were slight or negative; chronological and geographical attributions plateaued quicker than reconstructions; and there was early deterioration in reconstruction performance, highlighting the need for data augmentation techniques. As a result, three separate models were trained for each task and each corpus. Due to slower progress with the inscriptions dataset, characterized by more short entries, texts corresponding to less than 75 or more than 847 tokens were filtered out. Data augmentation techniques, as detailed in the method section above, were applied. Additionally, the decision to mask 3–20 characters instead of 1–20 resulted in long and rambling reconstructions when the test set required the reconstruction of 1–2 characters. Therefore, samples with 1–2 characters masked were also included in this phase of training.
For the reconstruction task, performance plateaued after 6 additional epochs (for a total of 9 epochs) for the inscriptions, achieving a CER of 22.3%, top-1 accuracy of 61.0%, and top-20 accuracy of 74.2%. Training on the papyrus dataset plateaued after 7 additional epochs (for a total of 11 epochs), with a CER of 16.3%, top-1 accuracy of 69.9%, and top-20 accuracy of 80.7%.
For the geographical attribution task, performance plateaued after 4 additional epochs (for a total of 7 epochs) for the inscriptions, achieving a top-1 accuracy of 75.0% and top-3 accuracy of 83.0%. Training on the papyrus dataset for geographical attribution plateaued after 5 additional epochs (for a total of 9 epochs), with a top-1 accuracy of 66.4% and top-3 accuracy of 79.9%.
For the dating task, performance plateaued after 1 additional epoch (for a total of 4 epochs) for the inscriptions, with an average deviation of 26.2 years and a median deviation of 1 year from the actual dating span. Training on the papyrus dataset for dating also plateaued after 1 additional epoch (for a total of 4 epochs), with an average deviation of 21.7 years and a median deviation of 0 years from the actual dating span.
3.1.2 Re-merging
In the final step, the effects of re-merging the fine-tuned models with the original base model were assessed. The TrIm, Elect Sign & Merge method (TIES-Merging), where parameters that changed only slightly during fine-tuning are trimmed ([11]), was applied, using the toolkit Mergekit (github.com/arcee-ai/mergekit). This process led to a deterioration of 10% in geographical attribution accuracy for inscriptions and 5% for papyri, along with an increase in the average error for dating by 4 years for inscriptions and 3 years for papyri.
4 Discussion
Table 1: Comparison of models for restoration, geographical attribution, and dating tasks. The results for Epigr_1_Llama3.1-8b-Inscruct are from this study, while results from Assael, Sommerschield et al. (2022) are included for comparison.
Restoration Region Date
Method CER (%) Top-1 (%) Top-20 (%) Top-1 (%) Top-3 (%) Years
Epigr_1_Llama3.1-8b-Inscruct 22.5 60.9 77.5 75.0 83.7 26.2
Assael et al. (2022) 26.3 61.8 78.3 70.8 82.1 29.3
Papy_1_Llama3.1-8b-Inscruct 16.3 71.3 85.0 66.5 80.0 21.7
Pavlopoulos et al. (2023) 54
Table 1 summarizes the results and allows for a comparison between the models developed in this study and the figures reported by Assael and colleagues [4]. Notably, the datasets used for testing Epigr_1_Llama3.1-8b-Instruct were smaller than those employed by Assael and colleagues [4], with 1,856 samples for dating and 2,870 for geographical attribution, compared to their dataset of 7,811 samples (though it is possible that not all of their entries in the test set included chronological or geographical data). A full, comparable replication of their study using a model like LLaMA would provide a more definitive assessment of which approach yields the best results. However, the primary aim of this experiment was not to maximize numerical performance. The goal was to determine whether a simple, instruct-based approach is competitive and warrants further exploration. The findings suggest that this is the case, and that additional experimentation with smaller, larger and more advanced similar base models, along with various approaches to dataset formatting, masking, training, and model-merging, would be worthwhile.
The key takeaway is that the tools developed here at least approximate the performance of the state-of-the-art model. Additionally, they offer a significant advantage by being applicable when scholars have an approximate sense of the number of missing characters but lack precise information on the number of missing word boundaries—an issue commonly encountered in the study of ancient textual artifacts. Furthermore, the papyri models introduced in this study are the first specifically designed for handling such material.
One legitimate concern that arises is whether some of the texts in the test datasets might have been included in the trillions of tokens used during the pretraining of LLaMA 3.1 8B. Although these texts were not part of the fine-tuning dataset, it is possible that the model had previously encountered this data. To address this concern, the models were tested on two unedited papyri from Uppsala University Library, with 100 differently masked versions of each text generated for ad-hoc evaluation.
  • Pap.Ups 106 (see image at urn.kb.se/resolve?urn=urn:nbn:se:alvin:portal:record-188165
    and text in Appendix 1) was processed with a character error rate (CER) of 17.2%, a top-1 accuracy of 62.0%, and a top-20 accuracy of 84.0%. The model plausibly assigns the document to Oxyrhynchus and accurately dates it to 71 AD (the fourth year of Vespasian, 71–72 AD).
  • Pap.Ups 18 (see image at urn.kb.se/resolve?urn=urn:nbn:se:alvin:portal:record-187926
    and text in Appendix 1) is more heavily damaged, resulting in lower model performance, but still provided useful results with a CER of 28.9%, a top-1 accuracy of 55.0%, and a top-20 accuracy of 71.0%. The model plausibly places it in Oxyrhynchus and dates the fragment to 304 AD, approximately 11–12 years from the correct date of around 292 AD (the eighth year of Diocletian and the seventh of Maximian are mentioned).
For the reconstruction tasks, however, there were only insignificant deteriorations of 0.1–0.2% in CER and top-1 accuracy, but a 3% increase in top-20 accuracy. Hence, the model’s overall restoration capacity interestingly improved when it was re-infused with the base model.
5 Conclusion
This experiment demonstrated the effectiveness of fine-tuning pretrained causal language models for dating, geographic attribution, and text restoration in ancient Greek inscriptions and documentary papyri using a simple chat template. While the results are preliminary, they suggest that this simple approach not only comes close to but, in some respects, surpasses the current state of the art. The simplicity and scalability of this method, particularly the ease of fine-tuning newer models on the same data without much modification, is a key advantage. There is significant potential for further improvement through more refined data cleaning, data augmentation and masking strategies, optimized training parameters, and experimentation with smaller or larger and more advanced base models as they become available in the future.
Although more specialized models may eventually outperform this simple method, instruct-tuned pretrained causal language models can offer an efficient and easily implementable moving baseline. Researchers developing machine learning methods for textual criticism could either aim to surpass this baseline or incorporate it as part of a more comprehensive solution. Rather than viewing different models as being in competition, a more productive approach may involve combining them. A system that integrates several models, supported by an adjudicator or ranker model, could potentially achieve the best results. This collaborative approach, taking advantage of the complementary strengths of various forms of AI, holds great promise for developing powerful tools to advance philological research.